# LLM-wrapper: Black-Box Semantic-Aware Adaptation of Vision-Language Models for Referring Expression Comprehension

### v2

@@ 4.1.3 The LLMs and their fine-tuning
Our main experiments are conducted using two different LLMs: Mixtral 8x7B Instruct (Jiang et al., 2024) (v0.1) and Llama 3 8B Instruct (AI@Meta, 2024) with Hugging Face’s implementation. We use Hugging Face’s supervised fine-tuning pipeline (SFT) (HuggingFace, 2024), that allows to implement the training choices discussed in Section 3. Specifically, for Llama 3 8B, we train 352M parameters, which is 4.20% of the original model size. For Mixtral 8x7B, we train 114M parameters, which is 0.24% of the original size. To further study the impact of the LLM scale, we also experiment with two additional families of models, Gemma 2 (Gemma, 2024) and GPT-Neo (Gao et al., 2021). The latter is a class of LLMs, based on a replication of the GPT-3 architecture, trained on the large curated Pile dataset (Gao et al., 2021), and which was not ‘instructed’ (Ouyang et al., 2022). We train LLM-wrapper with Adam (Kingma, 2014), with a batch-size of four, until convergence. We discuss in Section 4.4 the sample efficiency of LLM-wrapper. Unless stated otherwise, we use a learning rate of $10^{-5}$ and a rank of $r=128$ for LoRA. These hyper-parameters work consistently across three datasets, two LLMs and two VLMs, demonstrating the robustness of LLM-wrapper. We study the performance robustness of LLM-wrapper for different hyper-parameters in Section 4.4.
@@ 4.2 LLM-wrapper can adapt VLMs to the REC task
Table 3: Main results of LLM-wrapper on the REC task, in P@1$\uparrow$, on RefCOCO/+/g datasets. ‘(T)’ and ‘(B)’ stand for the ‘SwinT’ and ‘SwinB’ backbones respectively.
Model RefCOCOg RefCOCO RefCOCO+
Adaptation access VLM LLM val-umd test-umd val-unc test-unc val-unc test-unc
$\emptyset$ (zero-shot) GD(T) 60.09 59.32 50.69 50.94 51.65 51.79
Fine-tuning White-box GD(B) 78.51 77.99 83.86 84.12 73.46 73.46
LLM-wrapper Black-box GD(T) Mixtral 77.57 $\uparrow$17.5 77.05 $\uparrow$17.7 74.61$\uparrow$23.9 73.46$\uparrow$22.5 60.32$\uparrow$8.7 60.08$\uparrow$8.3
LLM-wrapper Black-box GD(T) Llama3 78.12 $\uparrow$18.0 77.36 $\uparrow$18.0 74.78$\uparrow$24.1 73.98$\uparrow$23.0 64.18$\uparrow$12.5 63.82$\uparrow$12.0
$\emptyset$ (zero-shot) GDrec(T) 67.61 68.37 51.82 52.12 53.28 53.16
Fine-tuning White-box GDrec(B) 80.19 79.85 83.84 84.21 73.68 73.58
LLM-wrapper Black-box GDrec(T) Mixtral 78.47 $\uparrow$10.9 77.92 $\uparrow$9.6 72.61$\uparrow$20.8 71.48$\uparrow$19.4 63.79$\uparrow$10.5 63.69$\uparrow$10.5
LLM-wrapper Black-box GDrec(T) Llama3 78.25 $\uparrow$10.6 78.01 $\uparrow$9.6 73.97$\uparrow$22.2 73.07$\uparrow$21.0 64.13$\uparrow$10.9 64.08$\uparrow$10.9
$\emptyset$ (zero-shot) Flo2 68.28 66.90 56.32 57.01 53.71 54.43
Fine-tuning White-box Flo2 90.32 91.06 93.08 93.42 88.20 88.49
LLM-wrapper Black-box Flo2 Mixtral 77.74 $\uparrow$9.5 76.98 $\uparrow$10.1 67.29$\uparrow$11.0 66.62$\uparrow$9.6 57.38$\uparrow$3.7 57.48$\uparrow$3.1
LLM-wrapper Black-box Flo2 Llama3 77.94 $\uparrow$9.7 77.15 $\uparrow$10.3 70.19$\uparrow$13.9 70.00$\uparrow$13.0 61.95$\uparrow$8.2 61.87$\uparrow$7.4
@@
In Table 3, we report the performances of VLMs on the REC task in three settings: zero-shot off-the-shelf VLM, after classic white-box fine-tuning, and after black-box adaptation with LLM-wrapper.
@@
Our first observation confirms that while VLMs demonstrate remarkable zero-shot performance on new tasks and datasets, their performance is still significantly lower than models specifically adapted to the given task (white-box fine-tuning). As shown in Table 3, for both GD (and its variant GDrec) and Flo2, there is a notable gap in REC performance, ranging from -11.5 to -37 P@1, between the zero-shot and fine-tuned versions.
@@
Interestingly, we observe that for all combinations of VLMs and LLMs, LLM-wrapper can adapt VLMs to the new REC task, despite the black-box setting. For instance, on RefCOCOg, LLM-wrapper brings improvements over the zero-shot models ranging from +9.5 to +18 P@1. Notably, while our proposed variant GDrec outperforms GD by large margins in a zero-shot setting (e.g., +9.05 P@1-test) due to better subject identification, both models perform similarly when adapted with LLM-wrapper. This shows that LLM-wrapper particularly improves VLMs that lack abilities useful for the task, such as subject identification in the case of REC for GD. On RefCOCO, LLM-wrapper improves again by significant margins zero-shot VLMs, e.g., +23.0 P@1 (test) for GD with Llama 3 8B. Given that textual queries in RefCOCO are very short (3.6 words on average), subject identification is relatively easy for these queries. Therefore, these gains indicate that LLM-wrapper greatly helps to disambiguate the object of interest from the distractors. On RefCOCO+, the improvement brought by LLM-wrapper is lower but still significant (between +3.1 and +12.5 P@1). This is expected as RefCOCO+ is designed to exclude location words, and most referring expressions are thus purely appearance-based descriptions (Qiao et al., 2021) where LLM-wrapper brings less value for adaptation.
@@
We present some qualitative results in Figure 2, where we show predictions from Flo2 before and after being adapted with LLM-wrapper, with Llama 3 8B. We observe that the adaptation with LLM-wrapper enables better spatial understanding (Fig. 2(a), Fig. 2(d)), subject identification (Fig. 2(b), Fig. 2(e)) and relational reasoning (Fig. 2(b), Fig. 2(c), Fig. 2(e)).
Finally, while LLM-wrapper is not intended to outperform classic white-box fine-tuning, it can approach its performance in some settings (e.g., with GD and GDrec on RefCOCOg, despite a smaller vision backbone for LLM-wrapper). In other setups, a gap remains, as fine-tuned models specifically train their vision encoder for the task. Overall, we demonstrate that LLM-wrapper significantly improves VLMs’ zero-shot performance in a black-box manner, consistently across all different VLMs and LLMs.
(a) “The front most cow to the right of other cows”
(b) “Person on the skateboard”
@@
(d) “The tie at the second from the left”
(e) “A boy wearing a grey shirt”
@@
(f) “White plane in the front”
l
Figure 2: Qualitative results of Flo2 outputs, before and after adaptation by LLM-wrapper, provided with the query in caption. Adapting Flo2 (in orange) with LLM-wrapper (using Llama 3 8B, and represented in blue) leads to improved reasoning and better bounding box selection.
@@ 4.3 Benefits of wrapping VLMs’ outputs with LLM-wrapper
$\emptyset$ (zero-shot VLM) Flo2 68.28 66.90
LLM-wrapper Flo2 77.94 77.15
@@
Ensembling VLMs. One advantage of LLM-wrapper is its free-text input format, which allows to adapt to various inputs. Indeed, we show that LLM-wrapper can learn to ensemble outputs from different VLMs as shown in Table 4. Specifically, we concatenate the predictions of the two VLMs in the prompt described in Section 3. The results show that when ensembling GDrec and Flo2, scores are boosted by +3.0 P@1 (val-umd) and +2.1 P@1 (test-umd) when compared to those obtained with the best-performing VLM adapted with LLM-wrapper, namely GDrec. This demonstrates that LLM-wrapper is capable of reasoning on multiple sources and leveraging the strengths of different models.
Indeed, we observe that Flo2 has a high precision, while GD and GDrec have lower precision but high recall. We show qualitatively in Figure 3 an example where LLM-wrapper leverages the complementarity of the two models. The figure displays predictions from Flo2 (e), GDrec (f), and LLM-wrapper ensembling predictions from both Flo2 and GDrec (d). We observe that while Flo2 fails to detect the target object as a possible candidate (b), the additional proposals from GDrec (c) enable LLM-wrapper to find the correct object (d) that each independent models missed.
@@
Adaptation VLM (fine-tuning) VLM (inference) P@1 - val $\uparrow$ P@1 - test $\uparrow$
$\emptyset$ (zero-shot VLM) $\emptyset$ GDrec 67.61 68.37
LLM-wrapper Flo2 GDrec 73.90  $\uparrow$6.3 73.45 $\uparrow$5.1
$\emptyset$ (zero-shot VLM) $\emptyset$ Flo2 68.28 66.90
LLM-wrapper GDrec Flo2 73.86  $\uparrow$5.6 73.03 $\uparrow$6.1
@@
Transferring a trained LLM-wrapper to a new VLM. Another advantage of using text as the input format is that LLM-wrapper does not rely on model-specific activation values, making it transferable from one VLM to another. For instance, LLM-wrapper can be fine-tuned on Flo2 and transferred to GD or GDrec. This is illustrated in Table 5, where, for instance, when fine-tuned on GDrec’s or Flo2’s outputs, transferring it at inference time to the other model’s outputs gives an increase from +5.6 to +6.3 P@1 over zero-shot VLMs. This shows that during fine-tuning, LLM-wrapper learns spatial and semantic notions that generalize to other models. This capacity of LLM-wrapper is particularly useful for private models, such as GD-1.5 (Ren et al., 2024), where creating the training set can be expensive – for instance, getting predictions for RefCOCOg train would cost $\approx$ \$1,600 (\$20 per 1,000 API calls). To illustrate this use case, we use a RefCOCOg val subset corresponding to 300 free API calls to GD-1.5. When LLM-wrapper is fine-tuned on GDrec’s outputs and applied on GD-1.5’s outputs for inference, results are boosted by a significant +29.0 P@1. This showcases how LLM-wrapper can be used on continuously updated versions of models without the need to re-train it with each model update.

### v3

@@ 4.1.3 The LLMs and their fine-tuning
Our main experiments are conducted using two different LLMs: Mixtral 8x7B Instruct (Jiang et al., 2024) (v0.1) and Llama 3 8B Instruct (AI@Meta, 2024) with Hugging Face’s implementation. We use Hugging Face’s supervised fine-tuning pipeline (SFT) (HuggingFace, 2024), that allows to implement the training choices discussed in Section 3. Specifically, for a same LoRA’s rank $r$, we train 352M parameters for Llama 3 8B (which is 4.20% of the model original size) and 114M parameters for Mixtral 8x7B (which is 0.24% of its original size). To further study the impact of the LLM scale, we also experiment with two additional families of models, Gemma 2 (Gemma, 2024) and GPT-Neo (Gao et al., 2021). The latter is a class of LLMs, based on a replication of the GPT-3 architecture, trained on the large curated Pile dataset (Gao et al., 2021), and which was not ‘instructed’ (Ouyang et al., 2022). We train LLM-wrapper with Adam (Kingma, 2014), with a batch-size of four, until convergence. We discuss in Section 4.4 the sample efficiency of LLM-wrapper. Unless stated otherwise, we use a learning rate of $10^{-5}$ and a rank of $r=128$ for LoRA. These hyper-parameters work consistently across four datasets, two LLMs and two VLMs, demonstrating the robustness of LLM-wrapper. We study the performance robustness of LLM-wrapper for different hyper-parameters in Section 4.4. We provide ablations quantifying the impact of trainable parameters count in Appendix C and statistics on invalid LLM outputs in Appendix D.
@@ 4.2 LLM-wrapper can adapt VLMs to the REC task
Table 3: Main results of LLM-wrapper on the REC task, in P@1$\uparrow$, on RefCOCO/+/g and Talk2Car datasets. ‘(T)’ and ‘(B)’ stand for the ‘SwinT’ and ‘SwinB’ backbones respectively.
Model RefCOCOg RefCOCO RefCOCO+ Talk2Car
Adaptation access VLM LLM val-umd test-umd val-unc test-unc val-unc test-unc val test
$\emptyset$ (zero-shot) GD(T) 60.09 59.32 50.69 50.94 51.65 51.79 55.37 58.44
Fine-tuning White-box GD(B) 78.51 77.99 83.86 84.12 73.46 73.46 N/A N/A
LLM-wrapper Black-box GD(T) Mixtral 77.57 $\uparrow$17.5 77.05 $\uparrow$17.7 74.61$\uparrow$23.9 73.46$\uparrow$22.5 60.32$\uparrow$8.7 60.08$\uparrow$8.3 64.75 $\uparrow$9.4 67.14 $\uparrow$8.7
LLM-wrapper Black-box GD(T) Llama3 78.12 $\uparrow$18.0 77.36 $\uparrow$18.0 74.78$\uparrow$24.1 73.98$\uparrow$23.0 64.18$\uparrow$12.5 63.82$\uparrow$12.0 65.95 $\uparrow$10.6 68.61 $\uparrow$10.2
$\emptyset$ (zero-shot) GDrec(T) 67.61 68.37 51.82 52.12 53.28 53.16 47.64 51.04
Fine-tuning White-box GDrec(B) 80.19 79.85 83.84 84.21 73.68 73.58 N/A N/A
LLM-wrapper Black-box GDrec(T) Mixtral 78.47 $\uparrow$10.9 77.92 $\uparrow$9.6 72.61$\uparrow$20.8 71.48$\uparrow$19.4 63.79$\uparrow$10.5 63.69$\uparrow$10.5 63.89 $\uparrow$16.3 66.78 $\uparrow$15.7
LLM-wrapper Black-box GDrec(T) Llama3 78.25 $\uparrow$10.6 78.01 $\uparrow$9.6 73.97$\uparrow$22.2 73.07$\uparrow$21.0 64.13$\uparrow$10.9 64.08$\uparrow$10.9 63.97 $\uparrow$16.3 66.45 $\uparrow$15.4
$\emptyset$ (zero-shot) Flo2 67.91 66.16 55.94 57.21 53.31 54.26 46.78 47.53
Fine-tuning White-box Flo2 90.32 91.02 93.07 93.42 88.19 88.49 N/A N/A
LLM-wrapper Black-box Flo2 Mixtral  78.96 $\uparrow$11.1 77.69 $\uparrow$11.5 68.85$\uparrow$12.9  68.21$\uparrow$11.0 57.58$\uparrow$4.3 58.26$\uparrow$4.0 61.65 $\uparrow$14.9 65.14 $\uparrow$17.6
LLM-wrapper Black-box Flo2 Llama3  78.76 $\uparrow$10.9 78.03 $\uparrow$11.9 71.74$\uparrow$15.8 71.91$\uparrow$14.7 62.63$\uparrow$9.3 62.73$\uparrow$8.5 61.74 $\uparrow$15.0 65.84 $\uparrow$18.3
@@
In Table 3, we report the performances of VLMs on the REC task in three settings: zero-shot off-the-shelf VLM, after classic white-box fine-tuning, and after black-box adaptation with LLM-wrapper. The results are obtained by adapting the VLM using exclusively the training data specific to each benchmark.
@@
Our first observation confirms that while VLMs demonstrate remarkable zero-shot performance on new tasks and datasets, their performance is still significantly lower than models specifically adapted to the given task (white-box fine-tuning). As shown in Table 3, for GD, GDrec and Flo2, there is a notable gap in performance, ranging from -11.5 to -37.1 P@1, between the zero-shot and fine-tuned versions.
@@
Interestingly, we observe that for all combinations of VLMs and LLMs, LLM-wrapper can adapt VLMs to the new REC task, despite the black-box setting. For instance, on RefCOCOg, LLM-wrapper brings improvements over the zero-shot models ranging from +9.6 to +18.0 P@1. Notably, while our proposed variant GDrec outperforms GD by large margins in a zero-shot setting (e.g., +9.1 P@1-test) due to better subject identification, both models perform similarly when adapted with LLM-wrapper. This shows that LLM-wrapper particularly improves VLMs that lack abilities useful for the task, such as subject identification for GD. On RefCOCO, LLM-wrapper improves again by significant margins zero-shot VLMs, e.g., +23.0 P@1 (test) for GD with Llama 3 8B. Given that textual queries in RefCOCO are very short (3.5 words on average), subject identification is relatively easy. Therefore, these gains indicate that LLM-wrapper greatly helps to disambiguate the object of interest from the distractors. On RefCOCO+, the improvement brought by LLM-wrapper is lower but still significant (between +4.0 and +12.5 P@1). This is expected as RefCOCO+ is designed to exclude location words, and most referring expressions are thus purely appearance-based descriptions (Qiao et al., 2021) where LLM-wrapper brings less value for adaptation. On the driving Talk2Car dataset, LLM-wrapper’s results remain consistent with those observed on the classic human-centric RefCOCO/+/g benchmarks, with score boosts provided for all VLM/LLM combinations, ranging from +8.7 up to +18.3 P@1.
@@
We present some qualitative results in Figure 2, where we show predictions from Flo2 before and after adaptation with LLM-wrapper, using Llama 3 8B. We observe that adaptation with LLM-wrapper enables better subject identification (Fig. 2(a), Fig. 2(d)), spatial understanding (Fig. 2(b), Fig. 2(e)), relational reasoning (Fig. 2(a), Fig. 2(c)) and avoids selecting a more visible object when the ground truth is comparatively small (Fig. 2(d), Fig. 2(f)). We display additional qualitative results, including failure cases, in Appendix E.
LLM-wrapper is not designed to outperform classic white-box fine-tuning but demonstrates competitive performance in some settings. For example, with GD and GDrec on RefCOCOg, LLM-wrapper achieves comparable results despite using a smaller vision backbone. Additionally, LLM-wrapper can complement white-box fine-tuning by adapting VLMs already optimized for the REC task. As shown in Appendix A, applying LLM-wrapper on top of fine-tuned models does not degrade performances and, in some cases, provides a slight boost. This highlights LLM-wrapper’s compatibility with state-of-the-art methods and its ability to enhance existing adaptations. Finally, in Appendix B, we extend our evaluation to the related task of Referring Expression Segmentation, leading to consistent score boosts.
(a) “Person on the skateboard”
@@
(d) “She said something about a big sign on a fence. Maybe that is her ! Pull over here by this person and we will find out”
(e) “Try to get in front of the car that passed us on the left. He is driving like a madman”
@@
(f) “My friend said she would be standing on the corner waiting for me, I think that might be her, will you stop there ?”
Figure 2: Qualitative results of Flo2 on RefCOCOg (first row) and Talk2Car (second row), before and after adaptation with LLM-wrapper, provided with queries as captions. Adapting Flo2 (in orange) with LLM-wrapper (using Llama 3 8B, in blue) leads to improved reasoning and box selection.
@@ 4.3 Benefits of wrapping VLMs’ outputs with LLM-wrapper
$\emptyset$ (zero-shot VLM) Flo2 67.91 66.16
@@
Ensembling VLMs. One advantage of LLM-wrapper is its free-text input format, which allows adapting to various inputs. Indeed, we show that LLM-wrapper can learn to ensemble outputs from different VLMs, as shown in Table 4. Specifically, we concatenate the predictions of the two VLMs in the prompt described in Section 3. The results show that when ensembling GDrec and Flo2, scores are boosted by +2.5 P@1 (val-umd) and +2.1 P@1 (test-umd) when compared to those obtained with the best-performing VLM adapted with LLM-wrapper, namely Flo2. This demonstrates that LLM-wrapper is capable of reasoning on multiple sources and leveraging the strengths of different models.
Indeed, we observe that Flo2 has a high precision, while GD and GDrec have lower precision but high recall. We show in Figure 3 a qualitative example where LLM-wrapper leverages the complementarity of the two models. The figure displays predictions from Flo2 (Fig. 3(e)), GDrec (Fig. 3(f)), and LLM-wrapper ensembling predictions from Flo2 and GDrec (Fig. 3(d)). We observe that while Flo2 fails to detect the target object as a possible candidate (Fig. 3(b)), the additional proposals from GDrec (Fig. 3(c)) enable LLM-wrapper to find the correct object (Fig. 3(a)) that each independent model missed.
@@
VLM VLM P@1 - val $\uparrow$ P@1 - val $\uparrow$ P@1 - test $\uparrow$
Adaptation (fine-tuning) (inference) (subset 300) (full) (full)
$\emptyset$ (zero-shot VLM) $\emptyset$ GDrec 66.00 67.61 68.37
LLM-wrapper Flo2 GDrec 74.00 $\uparrow$8.0 73.90  $\uparrow$6.3 73.45 $\uparrow$5.1
$\emptyset$ (zero-shot VLM) $\emptyset$ Flo2 71.67 67.91 66.16
LLM-wrapper GDrec Flo2 75.33 $\uparrow$3.7 73.86  $\uparrow$6.0 73.03 $\uparrow$6.9
@@
Transferring a trained LLM-wrapper to a new VLM. Another advantage of using text as the input format is that LLM-wrapper does not rely on model-specific activation values, making it transferable from one VLM to another. For instance, LLM-wrapper can be fine-tuned on Flo2 and transferred to GD or GDrec. This is illustrated in Table 5, where, for instance, when fine-tuned on GDrec’s or Flo2’s outputs, transferring it at inference time to the other model’s outputs gives an increase from +5.1 to +6.9 P@1 over zero-shot VLMs. This shows that during fine-tuning, LLM-wrapper learns spatial and semantic notions that generalize to other models. This capacity of LLM-wrapper is particularly useful for private models, such as GD-1.5 (Ren et al., 2024a), where creating the training set can be expensive – for instance, getting predictions for RefCOCOg train would cost $\approx$ \$1,600 (\$20 per 1,000 API calls). To illustrate this use case, we use a RefCOCOg val subset corresponding to 300 free API calls to GD-1.5. When LLM-wrapper is fine-tuned on GDrec’s outputs and applied on GD-1.5’s outputs for inference, results are boosted by a significant +29.0 P@1. To ensure that the val-subset reflects the full val set’s difficulty, we evaluate GDrec and Flo2 on both sets. The results are consistent, confirming that the val-subset is a good proxy for the full set and that the performance gains from LLM-wrapper on GD-1.5 are significant This showcases how LLM-wrapper can be used on private models or on continuously updated versions of models without the need to re-train it with each model update.
Transferring a trained LLM-wrapper to new datasets. LLM-wrapper demonstrates strong generalization across datasets, reducing the need for training data and resources on new target domains. To demonstrate this property, we adapt Flo2 with LLM-wrapper on RefCOCOg and evaluate it on RefCOCO and RefCOCO+. Results in Table 6 show substantial gains over zero-shot Flo2, with improvements up to +13.1 P@1 on RefCOCO and +7.7 P@1 on RefCOCO+. Although slightly below results from direct fine-tuning on target datasets, these improvements demonstrate LLM-wrapper’s ability to transfer knowledge effectively. To test generalization to more complex scenarios, we evaluate Flo2 adapted with LLM-wrapper using RefCOCOg on HC-RefLoCo, a benchmark with no training split and longer, more complex, referring expressions, with an average length of 84 words. LLM-wrapper achieves +18.9 and +19.1 P@1 improvements on the validation and test sets, respectively, compared to zero-shot Flo2. Interestingly, when white-box fine-tuned Flo2 is transferred to HC-RefLoCo, its performance boost over zero-shot Flo2 is less than half of that achieved by LLM-wrapper. These results highlight LLM-wrapper’s ability to handle in a zero-shot setting complex referring expressions despite being fine-tuned on ten times shorter expressions. More details on the impact of input complexity on performance are given in Sec. C.2. Qualitative examples on HC-RefLoCo are shown in Sec. E.1.
Table 6: LLM-wrapper performance on zero-shot dataset transfer. Results obtained with Llama 3 8B. ‘FT’ stands for ‘fine-tuning’. With , Flo2 is fine-tuned using ‘a collection of public supervised data on a wide range of downstream tasks’, including but not limited to RefCOCO/+/g (Xiao et al., 2024).
Adaptation VLM Fine-tuning Data Inference Data P@1 - val $\uparrow$ P@1 - test $\uparrow$
$\emptyset$ (zero-shot VLM) Flo2 HC-RefLoCo 48.04 47.39
White-box FT Flo2 RefCOCO/+/g HC-RefLoCo 56.75 $\uparrow$8.7 55.62 $\uparrow$8.2
LLM-wrapper Flo2 RefCOCOg HC-RefLoCo 66.93 $\uparrow$18.9 66.45 $\uparrow$19.1
$\emptyset$ (zero-shot VLM) Flo2 RefCOCO 55.94 57.21
LLM-wrapper Flo2 RefCOCO RefCOCO 71.74 $\uparrow$15.8 71.91 $\uparrow$14.7
LLM-wrapper Flo2 RefCOCOg RefCOCO 69.00 $\uparrow$13.1 68.88 $\uparrow$11.7
$\emptyset$ (zero-shot VLM) Flo2 RefCOCO+ 53.31 54.26
LLM-wrapper Flo2 RefCOCO+ RefCOCO+ 62.63 $\uparrow$9.3 62.73 $\uparrow$8.5
LLM-wrapper Flo2 RefCOCOg RefCOCO+ 61.00 $\uparrow$7.7 61.07 $\uparrow$6.8